Logbooks Lab Maintenance Evaporator_1 Evaporator_2 Laser cutter Target Production Test-Stand RH-ISAC RH-Cyclotron RH-Meson Hall RH-Beamlines RH-ARIEL
  Test-Stand  Not logged in ELOG logo
Message ID: 1048     Entry time: Thursday, April 17, 2025, 16:11
Author: Dave Neilson 
Target/No.#: C#5 
Source:  
Station:  
Module:  
Target Oven W.O#:  
Heat Shield W.O#: 45235 - 5 
Status:  
Subject: LPSIS C#5 conditioning 

Thursday, April 17, 2025, 15:46  target installed. now roughing

Thursday, April 17, 2025, 17:12 Turbo pumps on.

Tuesday, April 22, 2025, 15:45    Water cooling on. no effect on vacuum

Tuesday, April 22, 2025, 15:46  ramping TGHT TBHT to 580A 230A at .5A/30s

 Wednesday, April 23, 2025, 09:21 slowly increasing HV BIAS/EE/EZL to 30kV, 2.5kV, 13kV,  watching for sparks.

Wednesday, April 23, 2025, 10:46   first spark seen at 28kV, ezl 13kV.

Wednesday, April 23, 2025, 12:35   at HV BIAS >42kV. have seen 3 large sparks outside of the chamber (not due to the target assembly, that appears to be quite stable).  leaving at 41kV until appears sparks have ceased.

Wednesday, April 23, 2025, 13:59  No sparks at 41kV. while increasing to 42kV, 3 sparks occurred over ~75min. say 25min spark period at 41.5kV

 

Wednesday, April 23, 2025, 14:00 dropped HV to 30kV, EZL to 13kV. turning on optics for scans.   (controls for TMS Q2/Q5 must be looked at. it often must be reconfigured in canbus settings. sometimes multiple times to turn it on.)

FC0: 2.12e-8 A

Wire scanners on ISAC fiducials setting:

RPM0A: 2025-04-23 C#5 RPM0A_1nA,  2025-04-23 C#5 RPM0A_33nA

RPM0B: 2025-04-23 C#5 RPM0B_1nA,  2025-04-23 C#5 RPM0B_33nA

Wednesday, April 23, 2025, 14:12  running mass scan. - 2025-04-23 C#5 conditioning mass scan

Mass 40 RPM6, ISAC fiducials:

2025-04-23 C#5 RPM6A_33nA

2025-04-23 C#5 RPM6B_33nA

Wednesday, April 23, 2025, 14:50  decreasing HV to 0 0 0

Wednesday, April 23, 2025, 15:01 decreasing TGHT TBHT to 0 0 at 1A/30s

 Thursday, April 24, 2025, 08:16   turning off turbo pumps

Attachment 1: 2025-04-23_C#5_conditioning_report.csv  908 Bytes  Uploaded Wednesday, April 23, 2025, 17:02  | Show | Hide all
ELOG V2.9.2-2455